### v1

# Reconstrucion of oriented matroids from Varchenko-Gelfand algebras

Abstract.
The algebra of $R$-valued functions on the set of chambers of a real hyperplane arrangement is called the Varchenko–Gelfand (VG) algebra. This algebra carries a natural filtration by the degree with respect to Heaviside functions, giving rise to the associated graded VG algebra. When the coefficient ring $R$ is an integral domain of characteristic $2$, the graded VG algebra is known to be isomorphic to the Orlik–Solomon algebra. In this paper, we study VG algebras over coefficient rings of characteristic different from $2$, and investigate to what extent VG algebras determine the underlying oriented matroid structures.
Our main results concern hyperplane arrangements that are generic in codimension $2$. For such arrangements, if $R$ is an integral domain of characteristic not equal to $2$, then the oriented matroid can be recovered from both the filtered and the graded VG algebras.
We also formulate an algorithm that is expected to reconstruct oriented matroids from VG algebras in the case of general arrangements.
1. Introduction
Varchenko and Gelfand [19] introduced the algebra $\operatorname{\mathcal{VG}}(\mathcal{A})_{R}$ of $R$-valued functions on the set of chambers of a real central hyperplane arrangement $\mathcal{A}$. The VG algebra is generated by Heaviside functions, which define a degree filtration $\operatorname{\mathsf{F}}^{\bullet}\operatorname{\mathcal{VG}}(\mathcal{A})$. Passing to the associated graded algebra, one obtains $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})=\operatorname{Gr}_{\operatorname{\mathsf{F}}^{\bullet}}^{\bullet}\operatorname{\mathcal{VG}}(\mathcal{A})$.
The Hilbert series of $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$ coincides with the Poincaré polynomial of the complexified complement $M(\mathcal{A})$. Moreover, if $R=\mathbb{F}_{2}$, the graded VG algebra $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})_{\mathbb{F}_{2}}$ is isomorphic to the cohomology ring $H^{\bullet}(M(\mathcal{A}),\mathbb{F}_{2})$, which is also isomorphic to the Orlik-Solomon algebra $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})_{\mathbb{F}_{2}}$ defined from the intersection lattice $L(\mathcal{A})$. The relationship between VG algebras and spaces associated with $M(\mathcal{A})$ has been extensively studied [6, 7, 14, 16].
On the other hand, the arrangement $\mathcal{A}$ determines an oriented matroid. The oriented matroid is a very strong invariant: for example, the homeomorphism type of the complexified complement $M(\mathcal{A})$ is determined by the oriented matroid of $\mathcal{A}$ [4]. VG algebras are clearly oriented-matroid invariants. Indeed, if the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic (equivalent up to reorientation and permutation of the ground set, see §2 for details), then the corresponding VG algebras are also isomorphic.
It is therefore natural to ask whether VG algebras over coefficient rings with $\operatorname{char}R\neq 2$ contain more information than in the $\mathbb{F}_{2}$-case. We expect that VG algebras encode information as rich as that of oriented matroids. The most ambitious conjecture in this direction is the following.
Conjecture 1.1.
(4.3) Let $R$ be an integral domain with $\operatorname{char}R\neq 2$, and let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be real arrangements. If $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})_{R}$ as graded $R$-algebras, then the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic.
The main results of this paper establish reconstructibility under certain genericity assumptions on $\mathcal{A}$.
Theorem 1.2 (Theorem 3.6).
Let $R$ be an integral domain with $\operatorname{char}R\neq 2$, and let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be real arrangements. Assume that $\mathcal{A}_{1}$ is generic in codimension $2$ (see Definition 3.5). If $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})_{R}$ as filtered $R$-algebras, then the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic.
Theorem 1.3 (Theorem 5.3).
Let $R$ be an integral domain with $\operatorname{char}R\neq 2$, and let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be real arrangements. Assume that $\mathcal{A}_{1}$ is generic in codimension $2$. If $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})_{R}$ as graded $R$-algebras, then the oriented matroids of $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic.
Clearly Theorem 1.3 implies Theorem 1.2 immediately. However, the proofs rely on different techniques. The proof of Theorem 1.2 reconstructs the tope graph of the arrangement by means of primitive idempotent elements of $\operatorname{\mathcal{VG}}(\mathcal{A})$, while the proof of Theorem 1.3 reconstructs the signed circuits of the arrangement using square-zero elements of $\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$ and the relations among them. For this reason, we present the two results separately.
These results show that VG algebras contain sufficient information to recover oriented matroids under genericity conditions. We also point out that none of the following invariants determine the VG algebra: the intersection lattice $L(\mathcal{A})$, the Orlik-Solomon algebra $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})$, or the Poincaré polynomial of $M(\mathcal{A})$ (Example 3.10, Example 5.2, Example 5.4). These relationships are summarized in Figure 1. (See also Section 2.4 for details.)
$\ell=3$$\operatorname{char}R=2$3.2)NO3.4)NO5.2)NO5.2)5.2)NONONONONONONONO5.1) $\begin{subarray}{c}\text{Filtered $R$-algebra}\\ \operatorname{\mathcal{VG}}(\mathcal{A}_{1})\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Graded $R$-algebra}\\ \operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Graded $R$-module}\\ \operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Signed circuits}\\ \operatorname{\mathcal{C}}(\mathcal{A}_{1})\simeq\operatorname{\mathcal{C}}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Tope graph}\\ \operatorname{\mathcal{T}}(\mathcal{A}_{1})\simeq\operatorname{\mathcal{T}}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Lattice}\\ L(\mathcal{A}_{1})\simeq L(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Graded $R$-algebra}\\ \operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Graded $R$-module}\\ \operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Homeomorphic}\\ M(\mathcal{A}_{1})\simeq M(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Homotopic}\\ M(\mathcal{A}_{1})\simeq M(\mathcal{A}_{2})\end{subarray}$$\begin{subarray}{c}\text{Char. poly.}\\ \chi(\mathcal{A}_{1},t)=\chi(\mathcal{A}_{2},t)\end{subarray}$
Figure 1. Solid arrows are the known implications. Dashed arrows represent conditional implications. Dotted arrows labeled “NO” indicate that counterexamples are known. Blue arrows correspond to implications addressed in this paper.
Finally, we will formulate a conjectural reconstruction algorithm for oriented matroids from VG algebras (4.5, Remark 5.5).
The paper is organized as follows. In §2, we fix notations and recall basic facts about hyperplane arrangements, VG algebras, and oriented matroids. As background, we also review the known implications among these notions (Figure 1).
In §3, we reconstruct tope graphs from filtered VG algebras. The key idea is that the set of chambers can be canonically identified with the set of primitive idempotent elements in $\operatorname{\mathcal{VG}}(\mathcal{A})$. A crucial observation is that the set of Heaviside functions can be recovered as the set of idempotent elements in $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$ (Lemma 3.8), under the assumption that $\mathcal{A}$ is generic in codimension $2$. We also exhibit a pair of generic arrangements that share the same intersection lattice but have non-isomorphic VG algebras.
In §4, we analyze the set of idempotent elements in $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$, which we call generalized Heaviside functions (Theorem 4.2). Based on this, we formulate a non-deterministic algorithm that produces a graph from the filtered VG algebra. We conjecture that the algorithm always terminates with the tope graph. We illustrate the procedure for the $A_{3}$ arrangement in Example 4.6.
In §5, we discuss how to reconstruct oriented matroids from the graded VG algebra. First, we describe the set of square-zero elements in $\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$, which is closely related to generalized Heaviside functions. Then, again assuming that $\mathcal{A}$ is generic in codimension $2$, we show that the set of signed circuits can be recovered from the square-zero elements using the presentation of the graded VG algebra.
2. Preliminaries
Unless otherwise stated, $R$ denotes an integral domain, and $\mathcal{A}$ always denotes a real central hyperplane arrangement.
2.1. Hyperplane arrangements and oriented matroids
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a linear hyperplane arrangement in the real vector space $V=\mathbb{R}^{\ell}$. The complement $V\setminus\bigcup_{i=1}^{n}H_{i}$ decomposes into disjoint open cones, called chambers (or topes in the context of oriented matroids). We denote the set of all chambers by $\operatorname{ch}(\mathcal{A})$.
Choose a defining linear form $\alpha_{i}\in V^{*}$ for each hyperplane $H_{i}=\alpha_{i}^{-1}(0)$. The positive and negative half-spaces are denoted by $H_{i}^{+}=\alpha_{i}^{-1}(\mathbb{R}_{>0})$ and $H_{i}^{-}=\alpha_{i}^{-1}(\mathbb{R}_{<0})$, respectively. For chambers $C,C^{\prime}\in\operatorname{ch}(\mathcal{A})$, the set of separating hyperplanes is defined by
$\operatorname{Sep}(C,C^{\prime})=\{H\in\mathcal{A}\mid\text{$H$ separates $C$ and $C^{\prime}$}\}.$ (2.1)
The tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$ (or $\operatorname{\mathcal{T}}(\alpha_{1},\dots,\alpha_{n})$) of $\mathcal{A}$ is the finite graph $(\operatorname{ch}(\mathcal{A}),E)$ with vertex set $\operatorname{ch}(\mathcal{A})$ and edge set
$E=\{\{C,C^{\prime}\}\mid\#\operatorname{Sep}(C,C^{\prime})=1\}.$ (2.2)
The graph distance between two vertices $C$ and $C^{\prime}$ is given by $d(C,C^{\prime})=\#\operatorname{Sep}(C,C^{\prime})$.
The arrangement $\mathcal{A}$ (or more precisely, the tuple of defining linear forms $\bm{\alpha}=(\alpha_{1},\dots,\alpha_{n})\in(V^{*})^{n}$) determines an oriented matroid. One possible formulation is via the set of covectors
$\operatorname{\mathcal{V}^{*}}(\bm{\alpha})=\{(\operatorname{sgn}\alpha_{1}(v),\dots,\operatorname{sgn}\alpha_{n}(v))\in\{+,-,0\}^{n}\mid v\in V\},$ (2.3)
where $\operatorname{sgn}:\mathbb{R}\to\{\pm,0\}$ is the usual sign function, i.e., $\operatorname{sgn}(x)=+$ if $x>0$, $\operatorname{sgn}(x)=$- if $x<0$, and $\operatorname{sgn}(x)=0$ if $x=0$.
The sign vectors of the coefficients of linear dependence relations are called vectors, and the set of vectors is denoted by
$\operatorname{\mathcal{V}}(\bm{\alpha})=\{(\operatorname{sgn}\lambda_{1},\dots,\operatorname{sgn}\lambda_{n})\in\{+,-,0\}^{n}\mid\lambda_{1}\alpha_{1}+\dots+\lambda_{n}\alpha_{n}=0,(\lambda_{i}\in\mathbb{R})\}.$ (2.4)
For $\bm{\sigma}\in\operatorname{\mathcal{V}}(\bm{\alpha})$, define the support, positive support, and negative support by $\operatorname{Supp}(\bm{\sigma})=\{i\in[n]\mid\sigma_{i}\neq 0\}$, $\operatorname{Supp}^{+}(\bm{\sigma})=\{i\in[n]\mid\sigma_{i}=+\}$, and $\operatorname{Supp}^{-}(\bm{\sigma})=\{i\in[n]\mid\sigma_{i}=-\}$, respectively. A support-minimal dependent set is called a circuit. The set of signed circuits is defined as
[…]
For a signed circuit $\bm{\sigma}=(\sigma_{1},\dots,\sigma_{n})\in\operatorname{\mathcal{C}}(\bm{\alpha})$, one has
$\bigcap_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}H_{i}^{+}\cap\bigcap_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}H_{i}^{-}=\emptyset.$ (2.6)
Note that for each hyperplane $H_{i}$, there are two possible choices of defining linear form, $\alpha_{i}$ and $\alpha_{i}$. Changing the sing of $\alpha_{i}$ does not affect the hyperplane arrangement itself. We say that covectors $\operatorname{\mathcal{V}^{*}}(\alpha_{1},\dots,\alpha_{n})$ and $\operatorname{\mathcal{V}^{*}}(\beta_{1},\dots,\beta_{n})$ are reorientation equivalent if there exist $\varepsilon_{1},\dots,\varepsilon_{n}\in\{\pm 1\}$ such that
$\operatorname{\mathcal{V}^{*}}(\alpha_{1},\dots,\alpha_{n})=\operatorname{\mathcal{V}^{*}}(\varepsilon_{1}\beta_{1},\dots,\varepsilon_{n}\beta_{n}).$ (2.7)
Two oriented matroids are said to be isomorphic if they are equivalent under a reorientation and a permutation of the ground set $n]=\{1,\dots,n\}$.
Two arrangements $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ are isomorphic if their associated oriented matroids are isomorphic. In particular, isomorphic arrangements yield isomorphic tope graphs, and the converse also holds.
Proposition 2.1 ([3], Theorem 4.2.14).
Two oriented matroids are isomorphic if and only if their tope graphs are isomorphic.
2.2. Intersection lattice and Orlik-Solomon algebra
See [15] for details of this section.
The set of intersections is defined by
$L(\mathcal{A})=\left\{\left.\bigcap_{H\in\mathcal{B}}H\right|\mathcal{B}\subset\mathcal{A}\right\}$ (2.8)
with a partial order given by $X\leq Y\Longleftrightarrow X\supseteq Y$ ($X,Y\in L(\mathcal{A})$), which is called the intersection lattice. The poset is called the intersection lattice. The data of the intersection lattice $L(\mathcal{A})$ is equivalent to the underlying matroid of $(\alpha_{1},\dots,\alpha_{n})$. Recall that the Möbius function $\mu:L(\mathcal{A})\to\mathbb{Z}$ is defined by $\mu(V)=1$ and $\mu(X)=-\sum_{V\leq Y<X}\mu(Y)$ for $V<X$. The characteristic polynomial of $\mathcal{A}$ is then defined as $\chi(\mathcal{A},t)=\sum_{X\in L(\mathcal{A})}\mu(X)t^{\dim X}$. By the definition of Möbius function, if $\mathcal{A}\neq\emptyset$, we have $\chi(\mathcal{A},1)=0$, equivalently, $(t-1)|\chi(\mathcal{A},t)$.
Let us recall the definition of the Orlik-Solomon algebra. Let $E=Re_{1}\oplus\cdots\oplus Re_{n}$ be a free $R$-module of rank $n$, and let $\wedge E=\bigoplus_{k=0}^{n}E^{\wedge k}$ be the exterior algebra. Define a linear map $\partial:E^{\wedge k}\to E^{\wedge(k-1)}$ by
$\partial e_{i_{1},\dots,i_{k}}=\sum_{p=1}^{k}(-1)^{p-1}e_{i_{1}}\wedge\cdots\wedge\widehat{e_{i_{p}}}\wedge\cdots\wedge e_{i_{k}}.$ (2.9)
A subset $S=\{i_{1},\dots,i_{k}\}\subset[n]$ is called dependent if the intersection $H_{S}:=H_{i_{1}}\cap\cdots\cap H_{i_{k}}$ has codimension strictly smaller than $k$. The Orlik-Solomon algebra $\operatorname{\mathsf{OS}}^{\bullet}_{R}(\mathcal{A})$ is defined as the quotient algebra
$\operatorname{\mathsf{OS}}_{R}^{\bullet}(\mathcal{A})=\frac{\wedge E}{\langle\partial e_{S}\mid S\subset[n]\mbox{ is dependent}\rangle}.$ (2.10)
Denote by $M(\mathcal{A})=\mathbb{C}^{\ell}\setminus\bigcup_{H\in\mathcal{A}}H\otimes\mathbb{C}$ the complexified complement of $\mathcal{A}$. Orlik and Solomon proved that the cohomology ring $H^{\bullet}(M(\mathcal{A}),R)$ is isomorphic to $\operatorname{\mathsf{OS}}^{\bullet}_{R}(\mathcal{A})$ as $R$-algebras. When $R=\mathbb{R}$, considering de Rham presentation of $H^{\bullet}(M(\mathcal{A}),\mathbb{R})$, the isomorphism is realized by the map $e_{i}\mapsto d\log\alpha_{i}/2\pi\sqrt{-1}$. The rank $b_{k}=\operatorname{rank}\operatorname{\mathsf{OS}}^{k}(\mathcal{A})$ is equal to the $k$-th Betti number of the space $M(\mathcal{A}$). It is also known that the coefficients of the characteristic polynomial are the Betti numbers with alternating sign. More precisely,
$\chi(\mathcal{A},t)=t^{\ell}-b_{1}t^{\ell-1}+b_{2}t^{\ell-2}-\cdots+(-1)^{\ell}b_{\ell}.$ (2.11)
We also note that the sum of Betti numbers is equal to the number of chambers: $\operatorname{ch}(\mathcal{A})=(-1)^{\ell}\chi(\mathcal{A},-1)=\sum_{k=0}^{\ell}b_{k}$ [22].
2.3. Filtered and graded Varchenko-Gelfand algebras
The set of all $R$-valued functions on $\operatorname{ch}(\mathcal{A})$ is called the Varchenko-Gelfand (VG) algebra, denoted
$\operatorname{\mathcal{VG}}(\mathcal{A})_{R}=\{f:\operatorname{ch}(\mathcal{A})\longrightarrow R\}.$
We sometimes omit the subscript $R$. As an $R$-algebra, $\operatorname{\mathcal{VG}}(\mathcal{A})_{R}$ is isomorphic to the direct product $R^{\operatorname{ch}(\mathcal{A})}$. Recall that an element $f\in\operatorname{\mathcal{VG}}(\mathcal{A})$ is called idempotent if $f^{2}=f$, and primitive idempotent if it can not be written as the sum of two nonzero idempotent elements. For $C\in\operatorname{ch}(\mathcal{A})$, let $1_{C}:\operatorname{ch}(\mathcal{A})\to\mathbb{R}$ be the characteristic function, defined by $1_{C}(C^{\prime})=\delta_{C,C^{\prime}}$. Then the sets of idempotents and primitive idempotents are given by
$\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))=$ $\left\{\left.\sum_{C\in S}1_{C}\right|S\subset\operatorname{ch}(\mathcal{A})\right\}$ (2.12)
$\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))=$ $\{1_{C}\mid C\in\operatorname{ch}(\mathcal{A})\}.$
Thus the set of primitive idempotents is canonically identified with $\operatorname{ch}(\mathcal{A})$.
Definition 2.2.
For each hyperplane $H_{i}$, define $x_{i}^{+}:\operatorname{ch}(\mathcal{A})\to R$ by
$x_{i}^{+}(C)=\begin{cases}1,&(C\subset H_{i}^{+})\\ 0,&(C\not\subset H_{i}^{+}).\end{cases}$ (2.13)
Similarly, define $x_{i}^{-}$ with respect to $H_{i}^{-}$. These are called the Heaviside functions. They satisfy the relations
$x_{i}^{+}+x_{i}^{-}=1,\ \ \ \ \ (x_{i}^{\pm})^{2}=x_{i}^{\pm},\ \ \ \ \ x_{i}^{+}\cdot x_{i}^{-}=0.$ (2.14)
The VG algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$ is generated by Heaviside functions (see Proposition 2.4). This yields a natural filtration according to degree in the Heaviside functions.
Definition 2.3.
Define the subspace $\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A})\subset\operatorname{\mathcal{VG}}(\mathcal{A})$ by
[…]
We set $\operatorname{\mathsf{F}}^{-1}\operatorname{\mathcal{VG}}(\mathcal{A})=\{0\}$. Clearly, $\operatorname{\mathsf{F}}^{0}\operatorname{\mathcal{VG}}(\mathcal{A})=R$. The increasing filtration $0=\operatorname{\mathsf{F}}^{-1}\subset\operatorname{\mathsf{F}}^{0}\subset\operatorname{\mathsf{F}}^{1}\subset\operatorname{\mathsf{F}}^{2}\subset\cdots$ is called the VG filtration. Equipped with this filtration, $\operatorname{\mathcal{VG}}(\mathcal{A})$ is sometimes referred to as the filtered VG algebra.)
Since $\operatorname{\mathsf{F}}^{i}\cdot\operatorname{\mathsf{F}}^{j}\subset\operatorname{\mathsf{F}}^{i+j}$, we may define
$\operatorname{\mathsf{VG}}^{k}(\mathcal{A})=\operatorname{\mathsf{F}}^{k}/\operatorname{\mathsf{F}}^{k-1},\ \ \operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})=\bigoplus_{k\geq 0}\operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$ (2.16)
which we call the graded VG algebra.
Proposition 2.4 (Varchenko and Gelfand [19]).
  • (1)
    $\operatorname{\mathsf{F}}^{\ell}\operatorname{\mathcal{VG}}(\mathcal{A})=\operatorname{\mathcal{VG}}(\mathcal{A})$.
  • (2)
    $\operatorname{\mathsf{VG}}^{k}(\mathcal{A})\simeq\operatorname{\mathsf{OS}}^{k}(\mathcal{A})$ as $R$-modules. In particular, $\operatorname{rank}\operatorname{\mathsf{VG}}^{k}(\mathcal{A})=b_{k}$.
  • (3)
    If $R=\mathbb{F}_{2}$ (or more generally, if $\operatorname{char}R=2$), then $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})$ as $\mathbb{F}_{2}$-algebras.
Proposition 2.5 ([19]).
Let $\mathcal{I}_{\mathcal{A}}\subset R[e_{1},\dots,e_{n}]$ be the ideal generated by the following elements:
  • $e_{i}^{2}-e_{i}$, for $i=1,\dots,n$;
  • for each signed circuit $\bm{\sigma}\in\operatorname{\mathcal{C}}(\bm{\alpha})$,
    $\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}e_{i}\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}(e_{i}-1)-\prod\limits_{i\in\operatorname{Supp}^{+}(\bm{\sigma})}(e_{i}-1)\cdot\prod\limits_{i\in\operatorname{Supp}^{-}(\bm{\sigma})}e_{i}.$ (2.17)
Then the map $e_{i}\mapsto x_{i}^{+}$ gives an isomorphism of $R$-algebras $R[e_{1},\dots,e_{n}]/\mathcal{I}_{\mathcal{A}}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\operatorname{\mathcal{VG}}(\mathcal{A})$.
Let $\bm{\sigma}\in\operatorname{\mathcal{C}}(\bm{\alpha})$ be a signed circuit with $\operatorname{Supp}(\bm{\sigma})=\{i_{1},\dots,i_{k}\}$. Then there exists a unique (up to scalar) nontrivial linear relation
$\lambda_{i_{1}}\alpha_{i_{1}}+\dots+\lambda_{i_{k}}\alpha_{i_{k}}=0.$ (2.18)
The leading term of (2.17) is a degree k1k-1 homogeneous polynomial
$\sum_{p=1}^{k}\operatorname{sgn}(\lambda_{i_{p}})\cdot e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{p}\}},$ (2.19)
where $e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{k}\}}=e_{i_{1}}\cdots\widehat{e_{i_{p}}}\cdots e_{i_{k}}$. This relation also appears in the combinatorial analogue of the Orlik-Terao algebra [5].
Proposition 2.6 ([14], Theorem 5.9).
Let $I_{\mathcal{A}}\subset R[e_{1},\dots,e_{n}]$ be the ideal generated by:
  • $e_{i}^{2}$, for $i=1,\dots,n$;
  • for each signed circuit $\bm{\sigma}\in\operatorname{\mathcal{C}}(\bm{\alpha})$ with $\operatorname{Supp}(\bm{\sigma})=\{i_{1},\dots,i_{k}\}$ and linear relation (2.18), $\sum_{p=1}^{k}\operatorname{sgn}(\lambda_{i_{p}})\cdot e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{p}\}}$ (as in (2.19))
Then the map $e_{i}\mapsto\overline{x_{i}^{+}}$ induces an isomorphism of graded $R$-algebras $R[e_{1},\dots,e_{n}]/I_{\mathcal{A}}\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$.
Note that when $\operatorname{char}R=2$, relation (2.19) coincides with (2.9), which explains the isomorphism in Proposition 2.4 (3).
Remark 2.7.
Cordovil [5, Corollary 2.8] proved that NBC sets form a basis of $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$. It follows that for a signed circuit $\bm{\sigma}\in\operatorname{\mathcal{C}}(\bm{\alpha})$, the elements $e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{2}\}},\dots,e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{k}\}}$ can be part of a basis of $\operatorname{\mathsf{VG}}^{k-1}(\mathcal{A})$. Moreover:
  • $e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{1}\}},e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{2}\}},\dots,e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{k}\}}$ form a circuit in $\operatorname{\mathsf{VG}}^{k-1}(\mathcal{A})$.
  • The relation $\sum_{p=1}^{k}\operatorname{sgn}(\lambda_{i_{p}})\cdot e_{\operatorname{Supp}(\bm{\sigma})\setminus\{i_{p}\}}=0$ is unique up to scalar multiple.
For the inductive arguments, the following results are also useful. Fix $H\in\mathcal{A}$. We consider the associated arrangements: the deletion $\mathcal{A}^{\prime}:=\mathcal{A}\setminus\{H\}$ and the restriction $\mathcal{A}^{\prime\prime}=\mathcal{A}^{H}$, where $\mathcal{A}^{H}$ denotes the induced arrangement in the ambient space $H\simeq\mathbb{R}^{\ell-1}$. $\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})$ is naturally embedded into $\operatorname{\mathcal{VG}}(\mathcal{A})$ as a subalgebra. Moreover, $\operatorname{\mathcal{VG}}(\mathcal{A})$ and $\operatorname{\mathcal{VG}}(\mathcal{A}^{H})$ are related as follows. Let $C\in\operatorname{ch}(\mathcal{A}^{H})$ be a chamber of the restriction. Then there are exactly two chambers $C^{+}$ and $C^{-}$ of $\mathcal{A}$ such that
  • $C^{+}\subset H^{+}$ and $C^{-}\subset H^{-}$, and
  • $C$ is contained in the closures of both $C^{+}$ and $C^{-}$.
(Figure 2.) Using $C^{\pm}$, define linear maps $\rho^{\pm}:\operatorname{\mathcal{VG}}(\mathcal{A})\to\operatorname{\mathcal{VG}}(\mathcal{A}^{H}),f\mapsto\rho^{\pm}f$ by
$(\rho_{H}^{+}f)(C):=f(C^{+}),\ \ (\rho_{H}^{-}f)(C):=f(C^{-}),$ (2.20)
and set $\rho_{H}:=\rho_{H}^{+}-\rho_{H}^{-}$.
[…]
Figure 2. $C_{1},C_{2}\in\operatorname{ch}(\mathcal{A}^{H})$ and their lifts $C_{1}^{\pm},C_{2}^{\pm}\in\operatorname{ch}(\mathcal{A})$.
Note that $(\rho_{H}f)(C)=0$ if and only if $f(C^{+})=f(C^{-})$. Hence $\rho_{H}f=0$ if and only if $f\in\operatorname{\mathcal{VG}}(\mathcal{A})$ descends to $f\in\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})$. We therefore have a short exact sequence
$0\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathcal{VG}}(\mathcal{A}^{H}).$ (2.21)
Furthermore, Varchenko and Gelfand proved the following.
Proposition 2.8 ([19]).
For any $H$ and $0\leq k\leq\ell$, the sequence
$0\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{F}}^{k-1}\operatorname{\mathcal{VG}}(\mathcal{A}^{H})\longrightarrow 0$ (2.22)
is exact. Moreover, $f\in\operatorname{\mathcal{VG}}(\mathcal{A})$ lies in $\operatorname{\mathsf{F}}^{k}\operatorname{\mathcal{VG}}(\mathcal{A})$ if and only if $\rho_{H}(f)\in\operatorname{\mathsf{F}}^{k-1}\operatorname{\mathcal{VG}}(\mathcal{A}^{H})$ for every $H\in\mathcal{A}$.
By the snake lemma, (2.22) induces a short exact sequence of the graded version.
$0\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A}^{\prime})\longrightarrow\operatorname{\mathsf{VG}}^{k}(\mathcal{A})\stackrel{{\scriptstyle\rho_{H}}}{{\longrightarrow}}\operatorname{\mathsf{VG}}^{k-1}(\mathcal{A}^{H})\longrightarrow 0$ (2.23)
The following are also useful and frequently used.
Proposition 2.9.
Let $f\in\operatorname{\mathcal{VG}}(A)$. Then:
  • $f\in\operatorname{\mathsf{F}}^{0}\operatorname{\mathcal{VG}}(\mathcal{A})$ if and only if $\rho_{H}(f)=0$ for all $H\in\mathcal{A}$;
  • $f\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$ if and only if $\rho_{H}(f)$ is a constant function on $H$ for all $H\in\mathcal{A}$.
Remark 2.10.
The correspondence between chambers and elements in Orlik-Solomon algebras (Proposition 2.4, (2)) has appeared several times in recent works. In [21, 2], a basis of of $H^{\ell}(M(\mathcal{A}),\mathbb{Z})$ corresponding to chambers (a “chamber basis”) was constructed via the canonical isomorphism $C]\in H^{\operatorname{BM}}_{\ell}(M(\mathcal{A}),\mathbb{Z})\stackrel{{\scriptstyle\simeq}}{{\longrightarrow}}H^{\ell}(M(\mathcal{A}),\mathbb{Z})$. This basis is a dual to a homology basis constructed by Morse-theoretic method [20]. More recently, Eur and Lam [8] extended these ideas to oriented matroids. In their work, the compatibility between the map $\rho$ and the residue map in Orlik-Solomon algebras plays a crucial role.
2.4. Motivation and background
The isomorphism class of an Oriented matroid determines not only the VG algebra and the intersection lattice, but also the topological space $M(\mathcal{A})$ itself. The Salvetti complex [18] describes the homotopy type of $M(\mathcal{A})$ in terms of oriented matroids. Furthermore, Björner and Ziegler [18, 4] proved that the oriented matroid recovers the homeomorphism type of the complexified complement $M(\mathcal{A})$. The relationship between topology of $M(\mathcal{A})$ and the combinatorial structure of $L(\mathcal{A})$ is one of the central topics in the theory of hyperplane arrangements. Many (non/partial)implications are known among these topological and combinatorial objects. Some representative results are as follows:
  • $(1)$
    When $\ell=3$, the homeomorphism type of $M(\mathcal{A})$ determines $L(\mathcal{A})$ [13].
  • $(2)$
    There exist arrangements $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $M(\mathcal{A}_{1})$ and $M(\mathcal{A}_{2})$ are homotopy equivalent, but $L(\mathcal{A}_{1})$ and $L(\mathcal{A}_{2})$ are not isomorphic [9, Example 3.1]. Homotopy equivalence between $M(\mathcal{A}_{1})$ and $M(\mathcal{A}_{2})$ implies that their cohomology rings are isomorphic. Hence this example also shows that there exist $\mathcal{A}_{1},\mathcal{A}_{2}$ with $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})$ as graded algebras, while $L(\mathcal{A}_{1})\not\simeq L(\mathcal{A}_{2})$.
  • $(3)$
    The example in $(2)$ occurs in dimension $\ell=3$. Together with $(1)$, this implies that $M(\mathcal{A}_{1})$ and $M(\mathcal{A}_{2})$ are homotopy equivalent, but not homeomorphic.
  • $(4)$
    There exist arrangements $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $L(\mathcal{A}_{1})\simeq L(\mathcal{A}_{2})$ as posets, but $M(\mathcal{A}_{1})$ and $M(\mathcal{A}_{2})$ are not homotopy equivalent. Such a pair was first constructed over $\mathbb{C}$ by Rybnikov in [17], and later over $\mathbb{R}$ [12, 1].
  • $(5)$
    There exist $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})$ and $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})$ are isomorphic as graded $R$-module (equivalently, $\chi(\mathcal{A}_{1},t)=\chi(\mathcal{A}_{2},t)$), but not isomorphic as graded algebras. See [10, Example 3.1], [11, Example 4.9, 4.10].
These results are summarized in Figure 1. As noted in Proposition 2.4, when $\operatorname{char}R=2$, the graded VG algebra $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$ is isomorphic to $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A})$. One of the purposes of this paper is to understand the extent to which filtered and graded VG algebras reflect oriented matroid structures when $\operatorname{char}R\neq 2$, in contrast to the case $\operatorname{char}R=2$.
3. From filtered VG algebras to tope graphs
3.1. Recovering from the set of Heaviside functions
Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be central real hyperplane arrangements. For the moment, we do not assume that they have the same rank. However, we can easily recover both the rank and the number of hyperplanes from the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A}_{i})$, as follows.
Proposition 3.1.
Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be arrangements. If $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})$ (or $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$), then $\mathcal{A}_{1}=\#\mathcal{A}_{2}$ and the two arrangements have the same rank.
Proof.
By Proposition 2.4, $\operatorname{rank}\operatorname{\mathsf{VG}}^{1}(\mathcal{A}_{i})=b_{1}=\#\mathcal{A}_{i}$. The rank of $\mathcal{A}_{i}$ is equal to the largest integer $k>0$ such that $\operatorname{\mathsf{VG}}^{k}(\mathcal{A}_{i})\neq 0$. ∎
To recover the oriented matroid, it is enough (by Proposition 2.1) to recover the tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$. The vertex set of the tope graph is identified with $\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$. Our next task is to recover the adjacency relations.
Definition 3.2.
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement. Denote the set of all Heaviside functions by
$\mathcal{H}(\mathcal{A})=\{x_{i}^{\pm}\mid i=1,\dots,n\}.$ (3.1)
We call the hyperplane $H_{i}$ the support of the Heaviside function $x_{i}^{\pm}$. (See Definition 3.7 for the general notion of support.)
Note that $x,x^{\prime}\in\mathcal{H}$ have the same support if and only if either $x=x^{\prime}$ or $x+x^{\prime}=1$. The following observation is straightforward.
Proposition 3.3.
Let $\mathcal{A}$ and $\mathcal{H}(\mathcal{A})$ be as above. For two chambers $C,C^{\prime}\in\operatorname{ch}(\mathcal{A})$, we have
$\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2\cdot\#\operatorname{Sep}(C,C^{\prime}).$ (3.2)
Proof.
By definition of Heaviside functions, chambers $C$ and $C^{\prime}$ are separated by $H_{i}$ if and only if $x_{i}^{\pm}\cdot 1_{C}\neq x_{i}^{\pm}1_{C^{\prime}}$. The result follows immediately. ∎
By Proposition 3.3, $C$ and $C^{\prime}$ are adjacent if and only if $\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2$. Therefore, once the set of Heaviside functions $\mathcal{H}(\mathcal{A})\subset\operatorname{\mathcal{VG}}(\mathcal{A})$ is known, the adjacency relation on the vertex set $\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ can be recovered algebraically. More precisely, the edge set can be constructed as
$\{\{1_{C},1_{C}^{\prime}\}\subset\operatorname{PrimIdem}(\mathcal{A})\mid\#\{x\in\mathcal{H}(\mathcal{A})\mid x\cdot 1_{C}\neq x\cdot 1_{C^{\prime}}\}=2\}$ (3.3)
The main difficulty of the reconstruction problem is that, a priori, the set of Heaviside functions $\mathcal{H}(\mathcal{A})$ is not known. From the definitions, we have $\mathcal{H}(\mathcal{A})\subseteq\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$, $\mathcal{H}(\mathcal{A})\subseteq\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ and $1\notin\mathcal{H}(\mathcal{A})$. Hence,
$\mathcal{H}(\mathcal{A})\subseteq\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}.$ (3.4)
Since the right-hand-side of (3.4) is described purely in terms of the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$, if equality holds in (3.4) then the tope graph can be recovered from $\operatorname{\mathcal{VG}}(\mathcal{A})$. In general, however, the inclusion is strict, as the following example shows.
Example 3.4.
Let $\mathcal{A}=\{H_{1},H_{2},H_{3}\}$ be lines in $\mathbb{R}^{2}$ as in Figure 3 (upper left). Define $y^{\pm}\in\operatorname{\mathcal{VG}}(\mathcal{A})$ as in Figure 3 (upper right). Then,
$y^{+}=x_{1}^{+}+x_{2}^{+}+x_{3}^{+}-1.$ (3.5)
The left-hand side ($y^{+}$) is clearly contained in $\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$. The right-hand side lies in $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$. (We can also verify this fact by checking $\rho_{H_{i}}(y^{\pm})$ is a constant function for any $i=1,2,3$ (Proposition 2.8).) Thus $y\in\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}$, yet $y$ is not a Heaviside function. Hence $\mathcal{H}(\mathcal{A})\subsetneqq\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}$. (We will see in Theorem 4.2 that $\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}=\{x_{1}^{\pm},x_{2}^{\pm},x_{3}^{\pm},y^{\pm}\}$ .)
In this situation, we can not algebraically distinguish $x_{1}^{\pm},x_{2}^{\pm},x_{3}^{\pm}$ from $y^{\pm}$. Indeed, the map $x_{1}^{+}\mapsto y^{-},x_{2}^{+}\mapsto x_{2}^{+},x_{3}^{+}\mapsto x_{3}^{+}$ defines an automorphism of the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$. In this example, the automorphism group $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ (see Section 3.3), of filtered algebra is isomorphic to the wreath product $\{\pm 1\}\wr\mathfrak{S}_{3}\simeq\{\pm 1\}^{3}\rtimes\mathfrak{S}_{3}$, a group of order $48$, which is strictly larger than the automorphism group of the tope graph (a group of order $12$).
$H_{1}$$H_{2}$$H_{3}$$\operatorname{\mathcal{T}}(\mathcal{A})$$y^{+}:=$$y^{-}:=011011011$110110110==111111000++110001111++001111110-$111111$
Figure 3. An element $y\in\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ that is not a Heaviside function. (Arrows indicate positive sides).
3.2. Arrangements generic in codimension 22
Definition 3.5.
An arrangement $\mathcal{A}$ is called generic in codimension $2$ if $\mathcal{A}_{X}=2$ for all $X\in L(\mathcal{A})$ with $\operatorname{codim}X=2$.
Our first main result shows that the tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$ can be recovered from the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$ when $\mathcal{A}$ is generic in codimension $2$.
Theorem 3.6.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be real arrangements and suppose that $\mathcal{A}_{1}$ is generic in codimension $2$. If $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})_{R}\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})_{R}$ as filtered algebras, then the tope graphs $\operatorname{\mathcal{T}}(\mathcal{A}_{1})$ and $\operatorname{\mathcal{T}}(\mathcal{A}_{2})$ are isomorphic.
Before proving Theorem 3.6, we introduce the notion of support for elements of $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$. Recall that $y\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$ if and only if $\rho_{H_{i}}(y)$ is a constant function on $H_{i}$ for every $i=1,\dots,n$ (Proposition 2.8).
Definition 3.7.
For $y\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$, define the set $\operatorname{Supp}(y)\subset\mathcal{A}$ by
$\operatorname{Supp}(y)=\{H\in\mathcal{A}\mid\rho_{H}(y)\neq 0\}.$ (3.6)
For $y\in\operatorname{\mathsf{F}}^{1}$ the following are immediate:
  • $y\in\operatorname{\mathsf{F}}^{0}\operatorname{\mathcal{VG}}(\mathcal{A})$ if and only if $\operatorname{Supp}(y)=\emptyset$.
  • $y\in\mathcal{H}(\mathcal{A})$ if and only if $\operatorname{Supp}(y)=1$.
  • For the element $y^{\pm}$ in Example 3.4, we have $\operatorname{Supp}(y^{\pm})=\mathcal{A}$.
Now we come to the essential step in the proof of Theorem 3.6.
Lemma 3.8.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}$ be an arrangement generic in codimension $2$. Then $\mathcal{H}(\mathcal{A})=\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}$.
Proof of Lemma 3.8.
Let $y\in\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\}$. Since $y\neq 0,1$, we have $\operatorname{Supp}(y)>0$. Assume $\operatorname{Supp}(y)\neq 1$, i.e., $\operatorname{Supp}(y)\geq 2$, and choose distinct $H_{1},H_{2}\in\operatorname{Supp}(y)$. Let $p\in H_{1}\cap H_{2}$ be a generic point, i.e., a point in $H_{1}\cap H_{2}$ which is not contained in other hyperplanes. Since $\mathcal{A}$ is generic in codimension $2$, the only hyperplanes through $p$ are $H_{1}$ and $H_{2}$, so a neighborhood of $p$ is divided into four chambers $C_{1},\dots,C_{4}$ as in Figure 4.
[…]
Figure 4. For chambers $C_{1},\dots,C_{4}$ around a generic point $p\in H_{1}\cap H_{2}$.
Because $y$ is idempotent, each $y(C_{i})$ is either 0 or $1$. Since $\rho_{H_{1}}(y)$ is constant, we have
$\rho_{H_{1}}(y)=y(C_{1})-y(C_{4})=y(C_{2})-y(C_{3})\neq 0.$ (3.7)
Since $\operatorname{char}R\neq 2$, this forces either [^1] If $\operatorname{char}R=2$, there are two more possibilities. $y(C_{1})=y(C_{2})=1,y(C_{3})=y(C_{4})=0$ or $y(C_{1})=y(C_{2})=0,y(C_{3})=y(C_{4})=1$. In either case, $\rho_{H_{2}}(y)=0$, contradicting $H_{2}\in\operatorname{Supp}(y)$. Hence, $\operatorname{Supp}(y)=1$ and $y$ is a Heaviside functions. ∎
Proof of Theorem 3.6.
By Lemma 3.8,
$\#\mathcal{H}(\mathcal{A}_{1})$ $=\#\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}_{1}))\setminus\{0,1\}$ (3.8)
$=\#\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}_{2}))\setminus\{0,1\}$
$\geq\mathcal{H}(\mathcal{A}_{2}).$
By Proposition 3.1, $\mathcal{H}(\mathcal{A}_{1})=\#\mathcal{H}(\mathcal{A}_{2})$, so $\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}_{2}))\setminus\{0,1\}=\mathcal{H}(\mathcal{A}_{2})$. Therefore the adjacency relations in the tope graphs (given by Proposition 3.3) agrees, so $\operatorname{\mathcal{T}}(\mathcal{A}_{1})\simeq\operatorname{\mathcal{T}}(\mathcal{A}_{2})$. ∎
3.3. Automorphism groups
We first consider the group of algebra automorphisms $\operatorname{Aut}_{\operatorname{\mathsf{alg}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ (not necessarily preserving the filtration). Any algebra isomorphism $f:\operatorname{\mathcal{VG}}(\mathcal{A})\to\operatorname{\mathcal{VG}}(\mathcal{A})$ preserves $\operatorname{PrimIdem}(\mathcal{A})\simeq\operatorname{ch}(\mathcal{A})$, hence induces a permutation of chambers $\overline{f}:\operatorname{ch}(\mathcal{A})\stackrel{{\scriptstyle\simeq}}{{\to}}\operatorname{ch}(\mathcal{A})$. Conversely, any permutation $\overline{f}:\operatorname{ch}(\mathcal{A})\stackrel{{\scriptstyle\simeq}}{{\to}}\operatorname{ch}(\mathcal{A})$ induces an algebra isomorphism $f:\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\simeq}}{{\to}}\operatorname{\mathcal{VG}}(\mathcal{A})$. Thus
$\operatorname{Aut}_{\operatorname{\mathsf{alg}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\simeq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})),$ (3.9)
which is a group of order $\operatorname{ch}(\mathcal{A})!$.
We next compare the automorphism group of the filtered algebra $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ with graph automorphism group $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))$ and the permutation group of chambers $\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A}))$.
Theorem 3.9.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}$ be a real arrangement in $\mathbb{R}^{\ell}$.
  • $(1)$
    There are natural inclusions
    $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A})).$ (3.10)
    In particular, $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ is finite.
  • $(2)$
    If $\mathcal{A}$ is generic in codimension $2$, then
    $\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))=\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$ (3.11)
Proof.
$(1)$ Since $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{alg}}}(\operatorname{ch}(\mathcal{A}))$, the inclusion $\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\operatorname{\mathsf{set}}}(\operatorname{ch}(\mathcal{A}))$ was observed above.
If $f:\operatorname{\mathcal{T}}(\mathcal{A})\stackrel{{\scriptstyle\simeq}}{{\to}}\operatorname{\mathcal{T}}(\mathcal{A})$ is a graph automorphism, it induces a permutation of chambers and hence an algebra automorphism of $\operatorname{\mathcal{VG}}(\mathcal{A})$. It remains to show that $f$ preserves the filtration, equivalently, $f$ preserves Heaviside functions.
If $C_{0},C_{1}\in\operatorname{ch}(\mathcal{A})$ are adjacent and separated by a hyperplane $H$ with $C_{0}\subset H^{+}$ and $C_{1}\subset H^{-}$, then
$\{C\in\operatorname{ch}(\mathcal{A})\mid C\subset H^{+}\}$ $=\{C\in\operatorname{ch}(\mathcal{A})\mid d(C,C_{0})<d(C,C_{1})\}$ (3.12)
$\{C\in\operatorname{ch}(\mathcal{A})\mid C\subset H^{-}\}$ $=\{C\in\operatorname{ch}(\mathcal{A})\mid d(C,C_{0})>d(C,C_{1})\}.$
(Note that $d(C,C_{0})=d(C,C_{1})\pm 1$.) The right-hand sides of (3.12) are expressed purely in graph-theoretic terms, hence are preserved by graph automorphisms. Thus the Heaviside functions are preserved, so $f\in\operatorname{Aut}_{\mathsf{filt}}(\operatorname{\mathcal{VG}}(\mathcal{A}))$.
$(2)$ Let $f:\operatorname{\mathcal{VG}}(\mathcal{A})\stackrel{{\scriptstyle\simeq}}{{\to}}\operatorname{\mathcal{VG}}(\mathcal{A})$ be a filtered automorphism. Then $f$ preserves $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$ and $\operatorname{Idem}(\mathcal{A})$, and by Lemma 3.8, it preserves $\mathcal{H}(\mathcal{A})$. By Proposition 3.3, adjacency of the chambers is determined by $\mathcal{H}(\mathcal{A})$, so $f$ induces a graph automorphism. Hence $\operatorname{Aut}_{\mathsf{filt}}(\operatorname{\mathcal{VG}}(\mathcal{A}))\subseteq\operatorname{Aut}_{\mathsf{graph}}(\operatorname{\mathcal{T}}(\mathcal{A}))$, proving equality. ∎
Example 3.4 shows that the strict inclusion
$\operatorname{Aut}_{\operatorname{\mathsf{graph}}}(\operatorname{\mathcal{T}}(\mathcal{A}))\subsetneq\operatorname{Aut}_{\operatorname{\mathsf{filt}}}(\operatorname{\mathcal{VG}}(\mathcal{A})).$ (3.13)
can occur in general when $\mathcal{A}$ is not generic in codimension $2$.
3.4. The intersection lattice does not determine the filtered VG algebra
In this section, we exhibit an explicit pair of arrangements $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ such that $L(\mathcal{A}_{1})\simeq L(\mathcal{A}_{2})$ as lattices but $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})\not\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})$ as filtered algebras. Thus the intersection lattice does not determine the filtered Varchenko-Gelfand algebra.
Based on Theorem 3.6, it is enough to construct a pair of arrangements such that
  • they are generic in codimension $2$,
  • their intersection lattices are isomorphic,
  • their tope graphs are not isomorphic.
We can construct such a (well-known) pair using generic $6$-planes in $\mathbb{R}^{3}$.
Example 3.10.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be six generic planes in $\mathbb{R}^{3}$ as in Figure 5.
The intersection lattices $L(\mathcal{A}_{1})$ and $L(\mathcal{A}_{2})$ are isomorphic. However, tope graphs $\operatorname{\mathcal{T}}(\mathcal{A}_{1})$ and $\operatorname{\mathcal{T}}(\mathcal{A}_{2})$ are not isomorphic. Indeed, $\operatorname{\mathcal{T}}(\mathcal{A}_{1})$ has two vertices of degree $6$, whereas, $\operatorname{\mathcal{T}}(\mathcal{A}_{2})$ has none.
These arrangements are clearly generic in codimension $2$. Therefore, $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})\not\simeq\operatorname{\mathcal{VG}}(\mathcal{A}_{2})$ as filtered algebras.
$\mathcal{A}_{1}$$\mathcal{A}_{2}$
Figure 5. Six generic planes in $\mathbb{R}^{3}$.
4. Conjectural reconstruction algorithm
4.1. Generalized Heaviside functions
For simplicity, let
$\widetilde{\mathcal{H}}(\mathcal{A})=\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\operatorname{\mathcal{VG}}(\mathcal{A}))\setminus\{0,1\},$ (4.1)
and call an element $y\in\widetilde{\mathcal{H}}(\mathcal{A})$ a generalized Heaviside function. The key step in Section 3.2 (Lemma 3.8) was that if $\mathcal{A}$ is generic in codimension $2$, then
$\mathcal{H}(\mathcal{A})=\widetilde{\mathcal{H}}(\mathcal{A}).$ (4.2)
In this section, we analyze $\widetilde{\mathcal{H}}(\mathcal{A})$ more closely. The element in Example 3.4 is a generalized Heaviside function. We can generalize this construction as follows.
Example 4.1.
Let $H_{1},H_{2},\dots,H_{2s+1}$ be $2s+1$ ($s\in\mathbb{Z}_{\geq 0}$) hyperplanes such that $X=\bigcap_{i=1}^{2s+1}H_{i}$ is codimension $2$. We define the generalized Heaviside functions $\operatorname{Alt}^{\pm}(H_{1},\dots,H_{2s+1})$ as follows.
(Intuitive definition) There are $4s+2$ chambers. Assign values 0 and $1$ alternatingly to these chambers. This produces two idempotent elements $\operatorname{Alt}^{+}(H_{1},\dots,H_{2s+1})$ and $\operatorname{Alt}^{-}(H_{1},\dots,H_{2s+1})$ in $\widetilde{\mathcal{H}}(\mathcal{A})$ (Figure 6). Note that the case $s=0$ corresponds to an ordinary Heaviside function.
(More formal definition) Fix an orientation of the transversal plane to $X$. We may assume (after renumbering and reorienting if necessary) that $4s+2$ chambers $C_{1},C_{2},\dots,C_{4s+2}$ are arranged counterclockwise. Further assume $C_{1}=H_{1}^{+}\cap H_{2}^{+},C_{2}=H_{2}^{-}\cap H_{3}^{-},C_{3}=H_{3}^{+}\cap H_{4}^{+},\dots,C_{2s+1}=H_{2s+1}^{+}\cap H_{1}^{+},C_{2s+2}=H_{1}^{-}\cap H_{2}^{-},\dots$. Then, define
$\operatorname{Alt}^{+}(H_{1},\dots H_{2s+1})$= $1_{C_{1}}+1_{C_{3}}+1_{C_{5}}+\cdots+1_{C_{4s+1}}$ (4.3)
$=$ $x_{1}^{+}x_{2}^{+}+x_{3}^{+}x_{4}^{+}+\cdots x_{2s-1}^{+}x_{2s}^{+}+x_{2s+1}^{+}x_{1}^{+}$
$=$ $x_{1}^{+}+x_{2}^{+}+\cdots+x_{2s+1}^{+}-s,$
$\operatorname{Alt}^{-}(H_{1},\dots H_{2s+1})$= $1-\operatorname{Alt}^{+}(H_{1},\dots H_{2s+1}).$
$H_{1}$$H_{2}$$H_{3}$$H_{4}$$H_{5}$$C_{1}$$C_{2}$$C_{3}$$C_{4}$$C_{5}$$C_{6}$$C_{7}$$C_{8}$$C_{9}$$C_{10}$$\operatorname{Alt}^{+}(H_{1},\dots,H_{5})=$$\operatorname{Alt}^{-}(H_{1},\dots,H_{5})=011011011011011$
Figure 6. $\operatorname{Alt}^{+}(H_{1},\dots,H_{2s+1})$ and $\operatorname{Alt}^{-}(H_{1},\dots,H_{2s+1})$ ($s=2$).
Note that this construction works only when the number of hyperplanes $H_{1},\dots,H_{2s+1}$ is odd.
Before stating the main result, recall Sylvester-Gallai Theorem: Let $L_{1},\dots,L_{k}\subset\mathbb{RP}^{2}$ be lines in the real projective plane $\mathbb{RP}^{2}$. Then it has either the intersection $\bigcap_{i=1}^{k}L_{i}\neq\emptyset$ (pencil type), or there exists a point $p\in\mathbb{RP}^{2}$ through which exactly two of the lines pass.
The following theorem shows that every element in $\widetilde{\mathcal{H}}(\mathcal{A})$ is of the form $\operatorname{Alt}^{\pm}(H_{1},\dots,H_{2s+1})$.
Theorem 4.2.
Assume $\operatorname{char}R\neq 2$. Let $y\in\widetilde{\mathcal{H}}(\mathcal{A})$. Then there exist an odd number of hyperplanes $H_{1},\dots,H_{2s+1}\in\mathcal{A}$ ($s\geq 0$) such that the intersection $X=\bigcap_{i=1}^{2s+1}H_{i}$ has codimension $2$, and $y=\operatorname{Alt}^{\pm}(H_{1},\dots,H_{2s+1})$.
Proof.
Consider the support $\operatorname{Supp}(y)\subset\mathcal{A}$. If $\operatorname{Supp}(y)=1$, then $y$ is a Heaviside function. Thus we may assume $\operatorname{Supp}(y)\geq 2$.
If the intersection $\cap\operatorname{Supp}(y):=\bigcap_{H\in\operatorname{Supp}(y)}H$ has codimension $2$, then $\operatorname{Supp}(y)$ must be odd and $y$ is of the form $\operatorname{Alt}^{\pm}$.
Suppose instead that $\cap\operatorname{Supp}(y)$ has codimension at least $3$. Choose $X\in L(\operatorname{Supp}(y))$ with $\operatorname{codim}X=3$. Let $p\in X$ be a generic point and let $D\simeq D^{3}$ be a $3$-dimensional disk transversal to $X$ with $D\cap X=\{p\}$. Then the arrangement $\operatorname{Supp}(y)\cap D$ is isomorphic to the localization $\operatorname{Supp}(y)_{X}$ and $y|_{D}$ is a generalized Heaviside function of $\operatorname{Supp}(y)_{X}$.
By Sylvester-Gallai Theorem, there exist $H,H^{\prime}\in\operatorname{Supp}(y)_{X}$ such that no other hyperplanes in $\operatorname{Supp}(y)$ contain $H\cap H^{\prime}$. An argument similar to that in Lemma 3.8 then shows that this situation is impossible. Therefore, $\cap\operatorname{Supp}(y)$ has codimension at most $2$. ∎
4.2. Conjectural algorithm
In this section, let $R$ be an integral domain with $\operatorname{char}R\neq 2$. Even when the arrangement $\mathcal{A}$ is not generic in codimension $2$, the filtered algebra $\operatorname{\mathcal{VG}}(\mathcal{A})$ may still determine the tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$. We formulate the following conjecture.
Conjecture 4.3.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be real arrangements. If $\operatorname{\mathcal{VG}}(\mathcal{A}_{1})$ and $\operatorname{\mathcal{VG}}(\mathcal{A}_{2})$ are isomorphic as filtered algebras (or, if $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})$ and $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$ are isomorphic as graded algebras), then tope graphs $\operatorname{\mathcal{T}}(\mathcal{A}_{1})$ and $\operatorname{\mathcal{T}}(\mathcal{A}_{2})$ are isomorphic.
By Theorem 3.6, 4.3 holds for arrangements that are generic in codimension $2$ when $\operatorname{char}R\neq 2$. As seen in Section 4.1, however, there are many generalized Heaviside functions that are not Heaviside functions. To formulate a reconstruction procedure from $\operatorname{\mathcal{VG}}(\mathcal{A})$, we introduce the notion of “generalized tope graphs.”
Definition 4.4.
Let $y_{1},\dots,y_{p}\in\widetilde{\mathcal{H}}(\mathcal{A})$. The generalized tope graph $\operatorname{\widetilde{\mathcal{T}}}(y_{1},\dots,y_{p})$ is the graph defined as follows:
  • The vertex set is the set of primitive idempotents $\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$.
  • A pair $\{v,v^{\prime}\}\subset\operatorname{PrimIdem}(\operatorname{\mathcal{VG}}(\mathcal{A}))$ forms an edge if and only if
    $\{i\in[p]\mid y_{i}v\neq y_{i}v^{\prime}\}=1.$ (4.4)
Conjecture 4.5 (Non-deterministic reconstruction).
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement in $\mathbb{R}^{\ell}$. Perform the following steps:
  • (Step 1)
    Choose elements $y_{1},\dots,y_{n}\in\widetilde{\mathcal{H}}(\mathcal{A})$ such that $1,y_{1},\dots,y_{n}$ form a basis of $\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$.
  • (Step 2)
    If the generalized tope graph $\operatorname{\widetilde{\mathcal{T}}}(y_{1},\dots,y_{n})$ is the tope graph of some oriented matroid, stop. Otherwise, return to (Step 1) with a different choice of $y_{1},\dots,y_{n}$.
Then the resulting graph $\operatorname{\widetilde{\mathcal{T}}}(y_{1},\dots,y_{n})$ is isomorphic to the tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$.
Since $n=\#\mathcal{A}$ is determined purely from the filtered algebra structure (Proposition 3.1) and $\widetilde{\mathcal{H}}(\mathcal{A})$ is a finite set, the procedure above terminates in finitely many trials. Moreover, if the chosen $y_{1},\dots,y_{n}$ happen to be the Heaviside functions, then the graph obtained in (Step 2) is exactly the tope graph of $\mathcal{A}$, because condition (4.4) coincides with adjacency of chambers (Proposition 3.3).
It is worth noting that even if $y_{1},\dots,y_{n}$ are not (all) Heaviside functions, the graph $\operatorname{\widetilde{\mathcal{T}}}(y_{1},\dots,y_{n})$ may still be a tope graph of an oriented matroid (see the example below), and in many examples it is isomorphic to the original $\operatorname{\mathcal{T}}(\mathcal{A})$.
Example 4.6.
Let $\mathcal{A}=\{H_{1},H_{2},\dots,H_{6}\}$ be the $A_{3}$-arrangement in $\mathbb{R}^{3}$ defined by xyz(xy)(xz)(yz)=0xyz(x-y)(x-z)(y-z)=0. The characteristic polynomial is $\chi(\mathcal{A},t)=(t-1)(t-2)(t-3)$, and $\operatorname{ch}(\mathcal{A})=24$. We give numberings of chambers $1,2,\dots,12,1^{\prime},2^{\prime},\dots,12^{\prime}$ as in Figure 7.
$H_{2}$$H_{4}$$H_{6}$$H_{5}$$H_{3}$$H_{1}$$11^{\prime}22^{\prime}33^{\prime}44^{\prime}55^{\prime}66^{\prime}$778812121010991111
Figure 7. $\mathcal{A}=A_{3}$ and numbering of chambers (the chamber $i^{\prime}$ is the opposite chamber of the chamber ii). The remaining chambers are $7^{\prime},8^{\prime},9^{\prime},10^{\prime},11^{\prime},12^{\prime}$.
Let $x_{1},\dots,x_{6}$ be the Heaviside functions and define generalized Heaviside functions $y_{1},y_{2},y_{3},y_{4}$ as in Figure 8.
[…]
Figure 8. Heaviside functions $x_{1},\dots,x_{6}$ and generalized Heaviside functions $y_{1},\dots,y_{4}$.
For a chamber $C_{i}$, the product $x_{p}\cdot 1_{C_{i}}$ (or $y_{p}\cdot 1_{C_{i}}$) equals either $1_{C_{i}}$ or 0; the full product table appears in Figure 9. (where, we put $1$ if the product is $1_{C_{i}}$).
11 22 33 44 55 66 $1^{\prime}$ $2^{\prime}$ $3^{\prime}$ $4^{\prime}$ $5^{\prime}$ $6^{\prime}$ 77 88 99 $10$ $11$ $12$ $7^{\prime}$ $8^{\prime}$ $9^{\prime}$ $0^{\prime} 11$$1^{\prime}$ $12^{\prime}$
$x_{1}$ 0 0 0 0 0 0 11 11 11 11 11 11 0 0 0 0 0 0 11 11 11 11 11 11
$x_{2}$ 11 0 0 0 0 0 0 11 11 11 11 11 11 0 0 0 11 11 0 11 11 11 0 0
$x_{3}$ 11 11 0 0 0 0 0 0 11 11 11 11 11 11 11 11 11 11 0 0 0 0 0 0
$x_{4}$ 11 11 11 0 0 0 0 0 0 11 11 11 11 11 0 0 0 11 0 0 11 11 11 0
$x_{5}$ 11 11 11 11 0 0 0 0 0 0 11 11 0 0 0 0 0 0 11 11 11 11 11 11
$x_{6}$ 11 11 11 11 11 0 0 0 0 0 0 11 11 11 11 0 0 0 0 0 0 11 11 11
$y_{1}$ 11 0 11 11 11 11 0 11 0 0 0 0 11 0 0 0 11 11 0 11 11 11 0 0
$y_{2}$ 11 11 0 11 0 0 0 0 11 0 11 11 0 0 11 11 11 0 11 11 0 0 0 11
$y_{3}$ 0 0 0 0 11 0 11 11 11 11 0 11 11 11 11 0 0 0 0 0 0 11 11 11
$y_{4}$ 0 11 11 0 0 11 11 0 0 11 11 0 0 11 0 11 0 11 11 0 11 0 11 0
Figure 9. Product table between generalized Heaviside functions and $1_{C_{i}}$.
In Figure 10, we depict several generalized tope graphs. First, $\operatorname{\widetilde{\mathcal{T}}}(x_{1},x_{2},x_{3},x_{4},x_{5},x_{6})$ is the original tope graph $\operatorname{\mathcal{T}}(\mathcal{A})$. Although $y_{1},y_{2},y_{3}$ are not Heaviside functions, $\operatorname{\widetilde{\mathcal{T}}}(x_{1},x_{3},x_{5},y_{1},y_{2},y_{3})$ is a tope graph of an oriented matroid and is isomorphic to $\operatorname{\mathcal{T}}(\mathcal{A})$. On the other hand, $\operatorname{\widetilde{\mathcal{T}}}(y_{1},x_{2},x_{3},x_{4},x_{5},x_{6})$ and $\operatorname{\widetilde{\mathcal{T}}}(x_{1},x_{2},x_{4},y_{1},y_{2},y_{3})$ are not a tope graphs of oriented matroids, because they have vertices of degree $2$.
[…]
Figure 10. Some generalized tope graphs.
The point of 4.5 is that even when $y_{1},\dots,y_{n}$ are not Heaviside functions, if the resulting graph is an oriented matroid tope graph, then it must be isomorphic to $\operatorname{\mathcal{T}}(\mathcal{A})$.
5. From graded VG algebras to signed circuits
5.1. Square-zero elements
In the graded VG algebra, the square of a Heaviside class is zero. Thus, instead of idempotents, we consider the set of square-zero elements:
$\operatorname{\mathsf{N}^{2}}(\mathcal{A}):=\{\overline{y}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})\mid\overline{y}^{2}=0\mbox{ in }\operatorname{\mathsf{VG}}^{2}(\mathcal{A})\}.$ (5.1)
If $y\in\widetilde{\mathcal{H}}(\mathcal{A})\subset\operatorname{\mathsf{F}}^{1}\cap\operatorname{Idem}(\mathcal{A})$, then $y^{2}=y\in\operatorname{\mathsf{F}}^{1}$, hence $\overline{y}\in\operatorname{\mathsf{N}^{2}}(\mathcal{A})$. We now describe $\operatorname{\mathsf{N}^{2}}(\mathcal{A})$ in terms of generalized Heaviside functions.
Lemma 5.1.
Assume $\operatorname{char}R\neq 2$. Then,
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\bigcup_{y\in\widetilde{\mathcal{H}}(\mathcal{A})}R\cdot\overline{y}.$ (5.2)
Proof.
Since $\operatorname{\mathsf{VG}}^{1}=\operatorname{\mathsf{F}}^{1}/\operatorname{\mathsf{F}}^{0}$, the set of square-zero elements can be written as
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=\{y\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\mid y^{2}\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})\}/\operatorname{\mathsf{F}}^{0}.$ (5.3)
Let $y\in\operatorname{\mathsf{F}}^{1}$ and $y^{2}\in\operatorname{\mathsf{F}}^{1}$. Fix $H\in\operatorname{Supp}(y)$. By Proposition 2.9, $\rho_{H}(y)=\gamma_{1}$ is a nonzero constant function on $H$. Write $\operatorname{ch}(\mathcal{A}^{H})=\{D_{1},\dots,D_{s}\}$, and set $\alpha_{i}=y(D_{i}^{+})$, and $\beta_{i}=y(D_{i}^{-})$ (see Figure 11).
[…]
Figure 11. Square-zero elements.
Then,
$\alpha_{1}-\beta_{1}=\alpha_{2}-\beta_{2}=\cdots=\alpha_{s}-\beta_{s}=\gamma_{1}$ (5.4)
is a nonzero constant. Since $y^{2}\in\operatorname{\mathsf{F}}^{1}$, again by Proposition 2.9, there is a constant $\gamma_{2}$ satisfying
$\alpha_{1}^{2}-\beta_{1}^{2}=\alpha_{2}^{2}-\beta_{2}^{2}=\cdots=\alpha_{s}^{2}-\beta_{s}^{2}=\gamma_{2}.$ (5.5)
We also have
$\alpha_{1}+\beta_{1}=\alpha_{2}+\beta_{2}=\cdots=\alpha_{s}+\beta_{s}=\gamma_{3},$ (5.6)
where $\gamma_{3}=\gamma_{2}/\gamma_{1}$. Sinec $\operatorname{char}R\neq 2$, we have
$\alpha_{1}=\alpha_{2}=\cdots=\alpha_{s}=\frac{\gamma_{1}+\gamma_{3}}{2},\ \ \beta_{1}=\beta_{2}=\cdots=\beta_{s}=\frac{\gamma_{3}-\gamma_{1}}{2}.$ (5.7)
As in the proof of Theorem 4.2 and using the Sylvester-Gallai theorem, either
  • $(a)$
    $\operatorname{codim}\bigcap_{H\in\operatorname{Supp}(f)}H=2$, or
  • $(b)$
    there exist $H,H^{\prime}\in\operatorname{Supp}(y)$ such that no other hyperplane in $\operatorname{Supp}(y)$ contains $H\cap H^{\prime}$,
Case $(b)$ contradicts the fact that $H,H^{\prime}\in\operatorname{Supp}(f)$ (See Figure 4, by the above arguments, $y(C_{1})=y(C_{2})=y(C_{3})=y(C_{4})$ holds). Thus $(a)$ holds. It is easily seen that if $\operatorname{Supp}(f)$ must be odd; otherwise one can not assign the two constant values consistently. Hence $\operatorname{Supp}(f)=2s+1$ and the values alternate around the codimension $2$ intersection (Figure 12). Therefore, $\overline{y}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$ is a constant multiple of $\overline{\operatorname{Alt}^{+}(H_{1},\dots,H_{2s+1})}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$.
$H_{1}$$H_{2}$$H_{3}$$H_{4}$$H_{5}$$\alpha$$\beta$$\alpha$$\beta$$\alpha$$\beta$$\alpha$$\beta$$\alpha$$\beta$
Figure 12. $f=(\alpha-\beta)\cdot\operatorname{Alt}^{+}(H_{1},\dots,H_{2s+1})+\beta$.
By Theorem 4.2, we obtain (5.2). ∎
As an application of Lemma 5.1, we can distinguish non-isomorphic graded VG algebras even when the graded $R$-module structures agree.
Example 5.2.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be arrangements of six planes in $\mathbb{R}^{3}$ as in Figure 13. Their characteristic polynomials are $\chi(\mathcal{A}_{1},t)=\chi(\mathcal{A}_{2},t)=(t-1)(t-2)(t-3)$, so, $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})$ and $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$ are isomorphic as graded $R$-modules. It is known that $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\not\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})$ [10, Example 3.1]. Moreover, the graded VG algebras are also non-isomorphic $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\not\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$. Here we write $x_{i}$ for the class $\overline{x_{i}^{+}}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$. Then Lemma 5.1 gives
$\operatorname{\mathsf{N}^{2}}(\mathcal{A}_{1})=$ $\bigcup_{i=1}^{6}Rx_{i}\cup R(x_{1}+x_{2}-x_{3})\cup R(x_{1}-x_{4}+x_{5})$ (5.8)
$\cup R(x_{1}-x_{4}+x_{6})\cup R(x_{1}-x_{5}+x_{6})\cup R(x_{4}-x_{5}+x_{6})$
$\operatorname{\mathsf{N}^{2}}(\mathcal{A}_{2})=$ $\bigcup_{i=1}^{6}Rx_{i}\cup R(x_{1}+x_{2}-x_{3})\cup R(x_{1}+x_{5}-x_{6})$
$\cup R(x_{2}-x_{4}-x_{6})\cup R(x_{3}-x_{4}-x_{5}).$
The numbers of components are different, hence the graded VG algebras can not be isomorphic.
$H_{1}$$H_{6}$$H_{4}$$H_{5}$$H_{2}$$H_{3}$$\mathcal{A}_{1}$$H_{1}$$H_{5}$$H_{6}$$H_{3}$$H_{2}$$H_{4}$$\mathcal{A}_{2}$
Figure 13. $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ with the same characteristic polynomials
5.2. Recovering signed circuits
Theorem 5.3.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1},\mathcal{A}_{2}$ be real arrangements and suppose that $\mathcal{A}_{1}$ is generic in codimension $2$. If $\operatorname{\mathsf{VG}}(\mathcal{A}_{1})$ and $\operatorname{\mathsf{VG}}(\mathcal{A}_{2})$ are isomorphic as graded $R$-algebras, then the set of signed circuits $\operatorname{\mathcal{C}}(\mathcal{A}_{1})$ and $\operatorname{\mathcal{C}}(\mathcal{A}_{2})$ are isomorphic.
The idea of the proof is as follows. By the genericity assumption, the set of square-zero elements are exactly scalar multiples of Heaviside functions in $\operatorname{\mathsf{VG}}^{1}$. A refined version of Cordovil’s argument [5, Proposition 3.4] enables us to recover signed circuits.
Proof of Theorem 5.3.
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement in $\mathbb{R}^{\ell}$, and let $\alpha_{i}\in V^{*}$ be a defining equation of $H_{i}$. Assume $\mathcal{A}$ is generic in codimension $2$. We explain how to reconstruct signed circuits $\operatorname{\mathcal{C}}(\mathcal{A})$ from the graded VG algebra $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$.
By Lemma 3.8 we have $\widetilde{\mathcal{H}}(\mathcal{A})=\mathcal{H}(\mathcal{A})$, hence
$\operatorname{\mathsf{N}^{2}}(\mathcal{A})=R\cdot f_{1}\cup\cdots\cup R\cdot f_{n},$ (5.9)
for some $f_{1},\dots,f_{n}\in\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$. After permuting $[n]$ we may assume that
$R\cdot f_{i}=R\cdot\overline{x_{i}^{+}},$ (5.10)
as sets ($i=1,\dots,n$), where $x_{i}^{+}\in\operatorname{\mathsf{F}}^{1}\operatorname{\mathcal{VG}}(\mathcal{A})$ is the Heaviside function for $H_{i}$. Hence, there exists $\mu_{i}\in R^{\times}$ such that $f_{i}=\mu_{i}\cdot\overline{x_{i}^{+}}$ ($i=1,\dots,n$).
Let $I=\{i_{1},\dots,i_{k}\}\subset[n]$. By Cordovil [5, Corollary 2.5], the linear forms $\alpha_{i_{1}},\alpha_{i_{2}},\dots,\alpha_{i_{k}}$ are independent if and only if
$\overline{x_{i_{1}}^{+}}\cdot\overline{x_{i_{2}}^{+}}\cdots\overline{x_{i_{k}}^{+}}\neq 0,\mbox{ in }\ \operatorname{\mathsf{VG}}^{k}(\mathcal{A}),$ (5.11)
equivalent to $f_{i_{1}}\cdots f_{i_{k}}\neq 0$. (For simplicity, write $f_{I}=f_{i_{1}}\cdots f_{i_{k}}$.) Thus, $\{\alpha_{i_{1}},\dots,\alpha_{i_{k}}\}$ is a circuit if and only if
$f_{I}=0,\mbox{ and }f_{I\setminus\{i_{p}\}}\neq 0\mbox{ for $p=1,\dots,k$},$ (5.12)
where $I=\{i_{1},\dots,i_{k}\}$. Hence the matroid structure (equivalently the intersection lattice $L(\mathcal{A})$) of $\mathcal{A}$ is recovered from $f_{1},\dots,f_{n}$.
It remains to recover signed circuits. Let $I=\{i_{1},\dots,i_{k}\}\subset[n]$ be a circuit. Then there is a non trivial linear relation
$\lambda^{I}_{1}\cdot f_{I\setminus\{i_{1}\}}+\lambda^{I}_{2}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda^{I}_{k}\cdot f_{I\setminus\{i_{k}\}}=0,$ (5.13)
where $\lambda^{I}_{1},\dots,\lambda^{I}_{k}\in R$, and necessarily $\lambda^{I}_{p}\neq 0$ ($p=1,\dots,k$). Such a tuple $(\lambda_{1},\dots,\lambda_{k})$ is unique up to nonzero scalar multiple by an element in the quotient field of $R$.
Call generators $f_{1},\dots,f_{n}$ good if, for every circuit $I\subset[n]$, the relation (5.13) can be chosen so that $\lambda^{I}_{p}=\pm 1$ for all $p=1,\dots,k$. By Proposition 2.6, the classes $\overline{x_{1}^{+}},\dots,\overline{x_{n}^{+}}$ are good generators, therefore, good generators exist.
Fix good generators $f_{1},\dots,f_{n}$. For each circuit $I=\{i_{1},\dots,i_{k}\}\subset[n]$ choose a linear relation $\sum_{p=1}^{k}\sigma_{p}\cdot f_{I\setminus\{i_{p}\}}=0$ with $\sigma_{p}\in\{\pm 1\}$. Define the sign vector $(\lambda_{1},\dots,\lambda_{n})\in\{0,\pm\}^{n}$ by
$\lambda_{i}=\begin{cases}\sigma_{p},\ \mbox{$i\in I$ and $i=i_{p}$}\\ 0,\ i\notin I.\end{cases}$ (5.14)
We call $(\lambda_{1},\dots,\lambda_{n})$ the extension of $(\sigma_{p})$ by 0. Now, let us define $\check{\operatorname{\mathcal{C}}}(f_{1},\dots,f_{n})$ as
[…]
By Remark 2.7,
$\check{\operatorname{\mathcal{C}}}(\overline{x_{1}^{+}},\dots,\overline{x_{n}^{+}})=\operatorname{\mathcal{C}}(\alpha_{1},\dots,\alpha_{n}).$ (5.16)
It remains to show that if $(f_{1},\dots,f_{n})$ and $(f^{\prime}_{1},\dots,f^{\prime}_{n})$ are good generators with $R\cdot f_{i}=R\cdot f^{\prime}_{i}$ ($i=1,\dots,n$), then $\check{\operatorname{\mathcal{C}}}(f_{1},\dots,f_{n})$ and $\check{\operatorname{\mathcal{C}}}(f^{\prime}_{1},\dots,f^{\prime}_{n})$ differ by a reorientation.
Assume that $(f_{1},\dots,f_{n})$ and $(f^{\prime}_{1},\dots,f^{\prime}_{n})$ are good generators. For each circuit $I=\{i_{1},\dots,i_{k}\}\subset[n]$, we have relations
$\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+\lambda_{2}^{I}\cdot f_{I\setminus\{i_{2}\}}+\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}=0$ (5.17)
$\lambda_{1}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{1}\}}^{\prime}+\lambda_{2}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{2}\}}^{\prime}+\cdots+\lambda_{k}^{{}^{\prime}I}\cdot f_{I\setminus\{i_{k}\}}^{\prime}=0,$ (5.18)
with $\lambda_{p}^{I},\lambda_{p}^{{}^{\prime}I}\in\{\pm 1\}$. Since $R\cdot f_{i}=R\cdot f^{\prime}_{i}$, there exist $\mu_{i}\in R^{\times}$ such that $f_{i}=\mu_{i}\cdot f_{i}^{\prime}$ ($i=1,\dots,n$), hence
$\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I}\cdot f^{\prime}_{I\setminus\{i_{1}\}}+\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I}\cdot f^{\prime}_{I\setminus\{i_{2}\}}+\cdots+\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I}\cdot f^{\prime}_{I\setminus\{i_{k}\}}=0.$ (5.19)
Since (5.18) and (5.19) represent the same relation up to a unit, there exists $c_{I}\in R^{\times}$ such that
$(\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I},\mu_{I\setminus\{i_{2}\}}\lambda_{2}^{I},\dots,\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I})=c_{I}\cdot(\lambda_{1}^{{}^{\prime}I},\lambda_{2}^{{}^{\prime}I},\dots,\lambda_{k}^{{}^{\prime}I}).$ (5.20)
Comparing $p$-th and $q$-th component ($1\leq p<q\leq k$), we have $\mu_{I\setminus\{i_{p}\}}\lambda_{p}^{I}=c_{I}\lambda_{p}^{{}^{\prime}I}$ and $\mu_{I\setminus\{i_{q}\}}\lambda_{q}^{I}=c_{I}\lambda_{q}^{{}^{\prime}I}$, and therefore,
$\mu_{I\setminus\{i_{q}\}}=\frac{\lambda_{p}^{I}\cdot\lambda_{q}^{{}^{\prime}I}}{\lambda_{p}^{{}^{\prime}I}\cdot\lambda_{q}^{I}}\cdot\mu_{I\setminus\{i_{p}\}}=\pm\mu_{I\setminus\{i_{p}\}}.$ (5.21)
Since $\mu_{i_{p}}=\mu_{I}/\mu_{I\setminus\{i_{p}\}}$ and $\mu_{i_{q}}=\mu_{I}/\mu_{I\setminus\{i_{q}\}}$, we have $\mu_{i_{p}}=\pm\mu_{i_{q}}$. Thus, if indices $i,j\in[n]$ are contained in a single circuit $i,j\in I\subset[n]$, then $\mu_{i}=\pm\mu_{j}$. Let us define the equivalence relation $\sim$ on $[n]$ as the transitive closure of the binary relation
$i\sim j\Longleftrightarrow\mbox{there exists a circuit $I\subset[n]$ such that $i,j\in I$}.$ (5.22)
Then, clearly, $i\sim j$ implies $\mu_{i}=\pm\mu_{j}$. Let $n]=\bigsqcup_{\tau\in T}X_{\tau}$ be the partition into equivalence classes and choose a representative $r_{\tau}\in X_{\tau}$ from each equivalence class $X_{\tau}$. For each $i\in[n]$, there exists unique $\tau\in T$ such that $i\in X_{\tau}$. Now define $\widetilde{\mu}_{i}\in\{\pm 1\}$ by
$\widetilde{\mu}_{i}:=\frac{\mu_{i}}{\mu_{r_{\tau}}},\ \ (i\in X_{\tau}).$ (5.23)
Now let $\widetilde{f}_{i}:=\widetilde{\mu}_{i}\cdot f_{i}^{\prime}$. Then $\check{\operatorname{\mathcal{C}}}(f^{\prime}_{1},\dots,f^{\prime}_{n})$ and $\check{\operatorname{\mathcal{C}}}(\widetilde{f}_{1},\dots,\widetilde{f}_{n})$ are connected by the reorientation map $(\sigma_{1},\dots,\sigma_{n})\mapsto(\widetilde{\mu}_{1}\cdot\sigma_{1},\dots,\widetilde{\mu}_{n}\cdot\sigma_{n})$ of sign vectors. We will also see $\check{\operatorname{\mathcal{C}}}(\widetilde{f}_{1},\dots,\widetilde{f}_{n})=\check{\operatorname{\mathcal{C}}}(f_{1},\dots,f_{n})$. Indeed, for any circuit $I=\{i_{1},\dots,i_{k}\}\subset[n]$, there exists $\tau\in T$ such that $I\subset X_{\tau}$. And then
$\lambda_{1}^{I}\cdot f_{I\setminus\{i_{1}\}}+$ $\cdots+\lambda_{k}^{I}\cdot f_{I\setminus\{i_{k}\}}$ (5.24)
$=$ $\mu_{I\setminus\{i_{1}\}}\lambda_{1}^{I}\cdot f^{\prime}_{I\setminus\{i_{1}\}}+\cdots+\mu_{I\setminus\{i_{k}\}}\lambda_{k}^{I}\cdot f^{\prime}_{I\setminus\{i_{k}\}}$
$=$ $\mu_{r_{\tau}}^{k-1}\cdot\widetilde{\mu}_{I\setminus\{i_{1}\}}\cdot\lambda_{1}^{I}\cdot f^{\prime}_{I\setminus\{i_{1}\}}+\cdots+\mu_{r_{\tau}}^{k-1}\cdot\widetilde{\mu}_{I\setminus\{i_{k}\}}\cdot\lambda_{k}^{I}\cdot f^{\prime}_{I\setminus\{i_{k}\}}$
$=$ $\mu_{r_{\tau}}^{k-1}\cdot(\lambda_{1}^{I}\cdot\widetilde{f}_{I\setminus\{i_{1}\}}+\cdots+\lambda_{k}^{I}\cdot\widetilde{f}_{I\setminus\{i_{k}\}}).$
So, the linear relations among $f_{I\setminus\{i_{1}\}},\dots,f_{I\setminus\{i_{k}\}}$ and those of $\widetilde{f}_{I\setminus\{i_{1}\}},\dots,\widetilde{f}_{I\setminus\{i_{k}\}}$ are the same. Thus, we have $\check{\operatorname{\mathcal{C}}}(f_{1},\dots,f_{n})=\check{\operatorname{\mathcal{C}}}(\widetilde{f}_{1},\dots,\widetilde{f}_{n})$. ∎
Example 5.4.
Assume $\operatorname{char}R\neq 2$. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be the generic $6$-plane arrangements from Example 3.10. We have $L(\mathcal{A}_{1})\simeq L(\mathcal{A}_{2})$, hence $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})$ as graded $R$-algebras. However, graded VG algebras are not isomorphic, $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\not\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$. Indeed, if $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$, then by Theorem 5.3, the signed circuits would be equivalent (up to reorientation and relabeling), which would force the tope graphs to be isomorphic, contradicting Example 3.10.
Remark 5.5.
Let $\mathcal{A}=\{H_{1},\dots,H_{n}\}$ be a real arrangement in $\mathbb{R}^{\ell}$. In analogy with Section 4.2, we propose a conjectural algorithm recovering $\operatorname{\mathcal{C}}(\mathcal{A})$ from $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})$ as follows.
  • (Step 1)
    Choose $f_{1},\dots,f_{n}\in\operatorname{\mathsf{N}^{2}}(\mathcal{A})$ so that $f_{1},\dots,f_{n}$ form a good generator of $\operatorname{\mathsf{VG}}^{1}(\mathcal{A})$.
  • (Step 2)
    If $\check{\operatorname{\mathcal{C}}}(f_{1},\dots,f_{n})$ is the set of signed circuits of some oriented matroid, stop. Otherwise, return to (Step 1) with different choice of $f_{1},\dots,f_{n}$.
We conjecture that the resulting oriented matroid is equivalent to the original one. This procedure terminates in finitely many steps whenever there are only finitely many choices in (Step 1), e.g., the case $R^{\times}<\infty$. Unlike the filtered case (4.5), there can be infinitely many possibilities in (Step 1) when $R^{\times}=\infty$, so termination is not guaranteed in full generality.
5.3. Concluding remarks
The graded VG algebra is obviously determined by the filtered VG algebra. On the other hand, it remains open whether the filtered structure can be reconstructed from the graded one.
Question 5.6.
Let $R$ be an integral domain. Can one recover the filtered $R$-algebra $\operatorname{\mathcal{VG}}(\mathcal{A})_{R}$ from the graded $R$-algebra $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A})_{R}$ directly?
By Theorem 5.3, the answer is affirmative when $\operatorname{char}R\neq 2$ and $\mathcal{A}$ is generic in codimension $2$ via the reconstruction of the oriented matroid. It would be interesting to know whether if we can reconstruct the filtered VG algebra directly from the graded one.
In 4.3, we assumed $\operatorname{char}R\neq 2$. Here, we briefly discuss the case $\operatorname{char}R=2$. If $\operatorname{char}R=2$, we can not recover the oriented matroid from the graded VG algebra. The counterexample can be obtained again by six generic planes. Let $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ be six generic planes in $\mathbb{R}^{3}$ as in Example 3.10. Then $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{2})$ when $\operatorname{char}R=2$, because
  • since both are generic arrangements, we have $L(\mathcal{A}_{1})\simeq L(\mathcal{A}_{2})$;
  • it follows $\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{1})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{2})$ for any coefficient ring $R$;
  • by Proposition 2.4, when $\operatorname{char}R=2$, $\operatorname{\mathsf{VG}}^{\bullet}(\mathcal{A}_{i})\simeq\operatorname{\mathsf{OS}}^{\bullet}(\mathcal{A}_{i})$ ($i=1,2$).
Nevertheless, the oriented matroids are not isomorphic.
On the other hand, reconstructability from the filtered VG algebra with $\operatorname{char}R=2$ remains unclear.
Question 5.7.
Assume $\operatorname{char}R=2$. Can one recover the oriented matroid from $\operatorname{\mathcal{VG}}(\mathcal{A})_{R}$?
We expect that a deeper understanding of the relation between filtered and graded VG algebras will shed further light on oriented matroids and geometry of the complement $M(\mathcal{A})$.
Acknowledgements
The second author is supported by JSPS KAKENHI JP23H00081.